OpenStudy (anonymous):

Easy easy, a quick challenge question. x=x^4 list all possible x values

OpenStudy (phi):

write it as x^4 - x = 0 factor out x x( x^3 - 1) = 0 x=0 and x^3= 1 now find the 3 roots of 1. If you accept complex numbers, there are 3 roots

OpenStudy (anonymous):

yes complex numbers too

OpenStudy (anonymous):

what about x=x^5

OpenStudy (phi):

same idea x^5-x=0 x(x^4-1)=0 x=0 and x^4-1=0 x^4=1 now find the 4 roots of 1. 1 and -1 are two of them, i and -i are the other two